1) Given below are account balances for Charlie Company: Gross sales, $94,000 Sales returns and allowances,...

1)

Given below are account balances for Charlie Company:

Gross sales, $94,000

Sales returns and allowances, $4,000

Selling expenses, $12,000

Cost of goods sold, $52,000

Interest expense, $3,000

How much is the gross profit margin? (enter your percentage as a decimal rounded to two decimal places. Example - enter 46% as .46)

2)

Dorchester Company had the following balances at the end of 2018 and 2019 respectively:

Net Credit Sales - $870,000 for 2018 and $962,000 for 2019.

Accounts Receivable - $84,000 for 2018 and $115,000 for 2019.

Allowance for Doubtful Accounts - $4,000 for 2018 and 5,500 for 2019

Calculate the accounts receivable turnover ratio to one decimal place.
Answer #1

Answers

  • Question #1

A

Gross sales

$94,000

B

Sales return and allowances

$4,000

C = A-B

Net Sales

$90,000

D

Cost of Goods Sold

$52,000

E = C-D

Gross Profits

$38,000

F = (E/C) x 100

Gross Profit margin

42%

Answer

  • Question #2

2018

2019

Account receivables

$84,000

$115,000

Allowance account

($4,000)

($5,500)

A

Account receivables, net of allowance

$80,000

$109,500

B = (80000 + 109500)/2

Average account receivables, net

$94,750

C

Net Credit sales [2019]

$962,000

D = C/D

Account receivables turnover

10.2